OpenStudy (katielong):

mass defect: I'm a little stuck on clarifying how to calculate the mass defect. Here's an example I'm working on - Uranium 238/92 decays to an alpha 4/2 and Th 234/90, calculate the mass defect?

OpenStudy (irishboy123):

from my book, 238U92 has mass: 238.048608u from Wiki, 234Th90 mass = 234.043601u from book, alpha 4He2 mass = 4.002603u Th + alpha = 234.043601u + 4.002603u = 238.046204u defect = U - (Th + alpha) = 238.048608u - 238.046204u = 0.002404u convert to kgs or eV's (multiply by c^2). depends an awful lot on the accuracy of the masses i have used. can't find anything easily online.